Two angles are supplementary. The measure of the smaller angle is one- half of the measure of the greater angle. Find the measure of the greater angle.
Answer by amoresroy(333)

Let x = the measure of the smaller angle
180-x = the measure of the greater angle
x = (1/2)(180-x)
2x = 180 - x
3x = 180
x = 60
Greater angle = 180 - x = 180 - 60 = 120
The measure of the greater angle is 120 degrees.